OpenStudy (anonymous):

solve the equation by factoring 10x^2 + 31x + 15 = 0

OpenStudy (anonymous):

the choices are a. (-3/5, -5/2) b. (3/5, -5/2) c. (3/5, 5/2) d. (-3/10, -1/3)

OpenStudy (anonymous):

10x^2 +25x +6x +15 =0 5x(2x+5) +3(2x+5) = 0 (2x+5)(5x+3) = 0 x = (-3/5, -5/2)